Write the equation of the line through (-2, 3) and parallel to y = 2x - 4 in slope-intercept form (y = mx + b).

Answer by anne(47)

ok
so parallel lines are lines with opposite slope
the equation is y=2x-4
since opposite slopes the slope for the new line is -1/2
for y intercept
the two points we have are (-2,3)
so we substitute slope and the points in the equation to find the y intercept

y=mx+b
3= -1/2x+b
3= -1/2(-2) +b as here we reduce and 2 and 2 is gone.
3= 1+b
3-1= 1-1+b as here -1 and 1 reduce to zero
2= b
so our y intercept is 2



So our final equation is y= -1/2x+2


did this help?
CHEERS